OpenStudy (anonymous):

Solve this system of linear equations. -6x=26-20y -8x=8-20y

OpenStudy (bahrom7893):

-6x-26 = -20y or -20y = -6x-26 Plug that into the 2nd equation

OpenStudy (bahrom7893):

-8x = 8 -6x-26

OpenStudy (bahrom7893):

Solve that for x and then plug x into any of the y equations to get the answer.

OpenStudy (anonymous):

x=9 ? right ?

OpenStudy (bahrom7893):

let me check: -8x+6x=8-26 -2x = -18 yup x=9

OpenStudy (anonymous):

& y=4 ? :)

OpenStudy (bahrom7893):

-6*9-26 = -20y -54-26=-20y -80=-20y y = 4 YAY!

OpenStudy (anonymous):

Yayyyyyyy <3